Guliman Co is forecasting sales of 61 700 units of product for August. To make one unit of finished product,3 pounds o raw materials are required. Actual beginning and desired ending inventories of raw materials and finished goods are as follows: August 1 ugust 31 (Desired) 67.900 8,370 Actual) 73,100 Raw materials (pounds) Finished goods (units) Required: a. Calculate the number of units of product to be produced during Augusit. units be p b. Calculate the number of pounds of raw materials to be purchased during August.
Answer #1

a) Number of units be produced = 61700+8370-5740 = 64330 Units

b) Number of pounds purchased = (64330*3)+67900-73100 = 187790 pounds